# Improving Robustness to Model Inversion Attacks via Sparse Coding Architectures

### v1

@@ 4.1 Experimental results overview
[…]
Accuracy
CelebA
No-Defense
$11.35$
$0.718$
$256.4$
$0.779$
Gaussian-Noise
$10.39$
$0.604$
$264.8$
$0.644$
GAN
$10.15$
$0.613$
$289.7$
$0.635$
Titcombe et al.[71]
$10.18$
$0.636$
$304.1$
$0.654$
Gong et al.[22]++
$10.02$
$0.556$
$381.5$
$0.672$
Gong et al.[22]
$10.11$
$0.595$
$350.5$
$0.614$
Peng et al.[58]
$9.90$
$0.514$
$402.8$
$0.728$
Hayes et al.[25]
$9.92$
$0.556$
$383.7$
$0.621$
Wang et al.[78]
$9.85$
$0.527$
$402.8$
$0.742$
Sparse-Standard
$9.84$
$0.539$
$374.7$
$0.728$
SCA0.1
$\mathbf{9.79}$
$\mathbf{0.451}$
$\mathbf{391.9}$
$\mathbf{0.726}$
SCA0.25
$\mathbf{9.45}$
$\mathbf{0.442}$
$\mathbf{411.0}$
$\mathbf{0.739}$
SCA0.5
$\mathbf{9.40}$
$\mathbf{0.440}$
$\mathbf{412.6}$
$\mathbf{0.723}$
Medical
No-Defense
$22.04$
$0.396$
$196.1$
$0.998$
MNIST
Gaussian-Noise
$21.83$
$0.382$
$209.4$
$0.862$
GAN
$21.77$
$0.427$
$219.0$
$0.998$
Gong et al.[22]++
$21.50$
$0.359$
$273.1$
$0.894$
Titcombe et al.[71]
$21.68$
$0.360$
$286.3$
$0.899$
Gong et al.[22]
$21.75$
$0.477$
$249.1$
$0.770$
Peng et al.[58]
$21.82$
$0.381$
$268.3$
$0.927$
Hayes et al.[25]
$21.72$
$0.337$
$259.7$
$0.823$
Wang et al.[78]
$21.71$
$0.322$
$211.7$
$0.937$
Sparse-Standard
$20.97$
$0.086$
$239.3$
$0.907$
SCA0.1
$\mathbf{21.19}$
$\mathbf{0.057}$
$\mathbf{253.5}$
$\mathbf{0.888}$
SCA0.25
$\mathbf{21.17}$
$\mathbf{0.075}$
$\mathbf{280.1}$
$\mathbf{0.882}$
SCA0.5
$\mathbf{20.06}$
$\mathbf{0.055}$
$\mathbf{288.8}$
$\mathbf{0.881}$
Accuracy. SCA’s improvements in reconstruction metrics also do not come at the cost of accuracy. Tables 1, 2, & 3 report the most accurate of each defense’s multiple runs, and Appendix 0.E gives further analysis. SCA’s accuracy equals (within $0.003$) or outperforms all baselines on CelebA. On MedicalMNIST, SCA’s accuracy beats half of baselines and is worse by $0.045$ vs. the best baseline, even though we do not do extra optimization or even basic parameter tuning on SCA, as our goal is to focus on the performance of the essential sparse coding approach.
Basic Sparse-Standard outperforms SOTA baselines.
Interestingly, our basic Sparse-Standard baseline outperforms the best baselines’ PSNR on both CelebA and Medical MNIST, and also outperforms baselines’ SSIM on the latter. SCA0.5 then outperforms Sparse-Standard on all metrics, obtaining SSIM and FID that are better by factors of $1.2$ to $1.6$ and $1.1$ to $1.2$, respectively and also slightly better PSNR. Thus, while Sparse-Standard offers an inferior defense vs. SCA, it still offers a fast practical defense for less privacy-critical domains. The fact that a simplistic one-shot sparse coding net outperforms sophisticated recent defenses underscores the promise of sparse coding approaches.
Figure 4: Qualitative comparisons among actual and reconstructed images (end-to-end setting) under SCA and other defense baselines on the Medical MNIST dataset.
Dataset
Defense
[…]
Accuracy
CelebA
No-Defense
$16.26$
$0.262$
$201.8$
$0.773$
Gaussian-Noise
$16.08$
$0.262$
$220.4$
$0.638$
GAN
$13.55$
$0.133$
$199.6$
$0.668$
Titcombe et al.[71]
$15.13$
$0.191$
$197.7$
$0.695$
Gong et al.[22]++
$13.10$
$0.032$
$204.8$
$0.704$
Gong et al.[22]
$13.15$
$0.119$
$199.6$
$0.682$
Peng et al.[58]
$13.78$
$0.141$
$218.8$
$0.716$
Hayes et al.[25]
$14.10$
$0.004$
$199.0$
$0.664$
Wang et al.[78]
$13.63$
$0.0011$
$203.2$
$0.744$
Sparse-Standard
$13.09$
$0.002$
$222.1$
$0.749$
SCA0.1
$\mathbf{12.89}$
$\mathbf{0.004}$
$\mathbf{228.5}$
$\mathbf{0.748}$
SCA0.25
$\mathbf{12.73}$
$\mathbf{0.004}$
$\mathbf{218.8}$
$\mathbf{0.737}$
SCA0.5
$\mathbf{12.42}$
$\mathbf{0.001}$
$\mathbf{231.9}$
$\mathbf{0.741}$
Medical
No-Defense
$31.48$
$0.935$
$10.66$
$0.998$
MNIST
Gaussian-Noise
$30.46$
$0.920$
$12.23$
$0.862$
GAN
$27.34$
$0.480$
$33.77$
$0.998$
Gong et al.[22]++
$18.37$
$0.353$
$81.52$
$0.894$
Titcombe et al.[71]
$21.33$
$0.431$
$30.60$
$0.899$
Gong et al.[22]
$21.52$
$0.436$
$64.88$
$0.770$
Peng et al.[58]
$19.05$
$0.420$
$107.9$
$0.908$
Hayes et al.[25]
$18.48$
$0.007$
$150.9$
$0.824$
Wang et al.[78]
$20.48$
$0.549$
$30.01$
$0.946$
Sparse-Standard
$14.79$
$0.119$
$250.6$
$0.907$
SCA0.1
$\mathbf{13.43}$
$\mathbf{0.004}$
$\mathbf{352.1}$
$\mathbf{0.888}$
SCA0.25
$\mathbf{12.32}$
$\mathbf{0.004}$
$\mathbf{375.9}$
$\mathbf{0.882}$
SCA0.5
$\mathbf{12.04}$
$\mathbf{0.003}$
$\mathbf{369.9}$
$\mathbf{0.881}$
@@
Metric evaluations. Table 2 reports performance in the end-to-end network setting (lower rows=better defense performance). In this setting, SCA’s performance advantage widens. In terms of PSNR & FID, training data reconstructions under SCA0.1 are degraded by factors of $1.1$ to $3.3$ and $1.1$ to $11.7$ compared to Peng et al. [58] and Wang et al.[78], respectively (but slightly worse SSIM vs. Wang et al. on CelebA). On all metrics, SCA0.1 outperforms Hayes et al.[25] and Titcombe et al. [71] by factors of $1.1$ to $2.4$ and $1.1$ to $107.7$, respectively. SCA0.1 also outperforms the data augmentation defense of Gong et al. [22] by factors of $1.01$ to $109$. Increasing SCA’s $\lambda$ to $0.5$ causes it to outperform the same baselines on all metrics by factors of $1.2$ to $141$ [58], $1.2$ to $183$ [78], $1.2$ to $4.0$[25], $1.2$ to $191$ [71], and $1.1$ to $145.3$ [22], respectively.
@@
SCA’s performance is also as stable or more stable than baselines’ performance over multiple runs. For example, Fig. 4(a) plots the means & std. devs. of SCA & baselines’ per-run FID (the standard metric for faces) over multiple runs for CelebA faces in the Plug-&-Play setting, and Fig. 4(b) plots this for PSNR (the standard metric for grayscale images) on Medical MNIST in the end-to-end setting. SCA obtains better performance while also exhibiting stability of defense performance on par with the best baseline (and better stability than most baselines). See also Appendix 0.E.
@@
[…]
Accuracy
CelebA
No-Defense
$16.49$
$0.302$
$185.8$
$0.766$
Gaussian-Noise
$15.44$
$0.227$
$191.1$
$0.753$
GAN
$15.57$
$0.253$
$176.7$
$0.646$
Titcombe et al.[71]
$14.99$
$0.144$
$194.2$
$0.725$
Gong et al.[22]++
$15.06$
$0.038$
$190.5$
$0.756$
Gong et al.[22]
$15.65$
$0.044$
$185.8$
$0.653$
Peng et al.[58]
$16.23$
$0.211$
$198.6$
$0.717$
Hayes et al.[25]
$15.06$
$0.005$
$178.8$
$0.672$
Wang et al.[78]
$14.82$
$0.173$
$189.6$
$0.652$
Sparse-Standard
$15.39$
$0.009$
$187.0$
$0.746$
SCA0.1
$\mathbf{15.05}$
$\mathbf{0.005}$
$\mathbf{178.7}$
$\mathbf{0.745}$
SCA0.25
$\mathbf{14.76}$
$\mathbf{0.003}$
$\mathbf{191.1}$
$\mathbf{0.743}$
SCA0.5
$\mathbf{14.71}$
$\mathbf{0.003}$
$\mathbf{206.1}$
$\mathbf{0.739}$
Medical
No-Defense
$23.47$
$0.776$
$45.57$
$0.993$
MNIST
Gaussian-Noise
$21.93$
$0.722$
$44.72$
$0.811$
GAN
$21.67$
$0.719$
$48.49$
$0.912$
Gong et al.[22]++
$21.07$
$0.573$
$67.53$
$0.931$
Titcombe et al.[71]
$21.35$
$0.704$
$48.82$
$0.961$
Gong et al.[22]
$21.33$
$0.720$
$41.74$
$0.925$
Peng et al.[58]
$18.98$
$0.426$
$124.8$
$0.914$
Hayes et al.[25]
$21.46$
$0.442$
$137.4$
$0.850$
Wang et al.[78]
$20.03$
$0.538$
$65.17$
$0.986$
Sparse-Standard
$15.33$
$0.149$
$142.4$
$0.955$
SCA0.1
$\mathbf{13.95}$
$\mathbf{0.008}$
$\mathbf{244.9}$
$\mathbf{0.946}$
SCA0.25
$\mathbf{12.31}$
$\mathbf{0.008}$
$\mathbf{255.3}$
$\mathbf{0.928}$
SCA0.5
$\mathbf{12.27}$
$\mathbf{0.001}$
$\mathbf{285.3}$
$\mathbf{0.909}$
@@
We also try varying SCA’s and Sparse-Standard’s sparsity parameters $\lambda$ and recompute PSNR, SSIM, FID, and accuracy. Table 6 in Appendix 0.A.6 shows that for each $\lambda$ and defense metric, SCA significantly outperforms the off-the-shelf Sparse-Standard architecture for a small accuracy cost. Thus, for a given $\lambda$ with Sparse-Standard, we can use a (smaller) $\lambda$ with SCA to obtain better reconstruction and higher or equal (within $0.0017$) accuracy. SCA is also amenable to more sophisticated tuning (and performance improvements) by tuning different $\lambda$ per sparse layer (e.g., by having a sparser representation of inputs but less sparse reductions of downstream layers). We avoid such tuning here as it is unnecessary for good performance.
@@ 5 Empirical analysis of sparse coding robustness to attack
Sparse-coding layers’ robustness to privacy attacks can be observed empirically. Consider that in both end-to-end and split network threat models, the attacker trains the attack to map leaked raw hidden layer outputs back to input images. Attacks are thus highly dependent on these outputs’ distributions. To visualize these distributions, recall that UMAP projections compute a 2D visualization of the global structure of distances between different training images’ features according to a particular layer [47]. Fig. 6 plots UMAP 2D projections of linear layer feature distributions of each training data input after either two linear layers (Figs. 5(a) & 5(d)), two convolutional layers (Figs. 5(b) & 5(e)), or two sparse coding layers (with interspersed dense layers – Figs. 5(c) & 5(f)).
Importantly, observe that after two linear or two convolutional layers, points are clustered by color, meaning that input images’ features are highly clustered by label (e.g. in MNIST nearly all $4$’s have similar features). This class-clustered property leaves such layers vulnerable to model inversion attacks, as an attacker can ‘home in on’ examples from a specific class. In contrast, the goal in sparse coding is not to optimize the classification objective by separating classes, but rather to jettison unnecessary information. Here, this means that unnecessary information is jettisoned both from the input image and also the downstream dense layer. Per Figs. 5(c) & 5(f), this tends to ‘uncluster’ remaining non-sparsified features of training examples from the same class, making it much harder for an attacker to compute informative gradients to home in on a training example.
@@ 6 Discussion & Conclusion
In this paper, we have provided the first study of neural network architectures that are robust to model inversion attacks. We have shown that a standard off-the-shelf sparse-coding architecture can outperform state-of-the-art defenses, and we have refined this idea to design an architecture that obtains superior performance. More broadly, we have shown that the natural properties of sparse coded layers can control the extraneous private information about the training data that is encoded in a network without resorting to complex and computationally intensive parameter tuning techniques. Our work reveals a deep connection between state-of-the-art privacy vulnerabilities and three decades of computer science research on sparse coding for other application domains.
Currently, our basic research implementation of SCA achieves compute times that in the worst-case are no better than some state-of-the-art baselines (see Table 15 in Appendix 0.F), and like the most performant benchmarks, SCA currently has no known provable privacy guarantees. However, given the rich theoretic body of work on fast algorithms and provable guarantees for sparse coding, we believe these aspects are opportune areas for improvement in future work.
@@ 0.A.1 Reproducibility
In order to promote further research and standardize the evaluations of new defenses, we provide full cluster-ready PyTorch [57] implementations of SCA and all benchmarks as well as replication codes for all experiments on our (author-anonymized) repository at: https://github.com/SayantonDibbo/SCA.
@@ 0.A.2 Adapting Rozell LCA to Convolutional Networks
Although the original LCA formulation [60] was introduced for the non-convolutional case, it is based on the general principle of feature-similarity-based competition between neurons within the same layer, which can be adapted to the convolutional setting via only two minimal changes to Equation 2 [70, 40]. In Rozell’s original formulation, $\Psi(t)$ can simply be recast from a matrix multiplication to a convolution between the input and dictionary. Second, the lateral interaction tensor, $\mathcal{G}$ in Equation 2, can also be recast from a matrix multiplication to a convolution between the dictionary and its transpose.

### v2

@@ 4.1 Experimental results overview
CelebA No-Defense $16.26$ $0.262$ $201.8$ $0.773$
Gaussian-Noise $16.08$ $0.262$ $220.4$ $0.638$
GAN $13.55$ $0.133$ $199.6$ $0.668$
Titcombe et al.[75] $15.13$ $0.191$ $197.7$ $0.695$
Gong et al.[25]++ $13.10$ $0.032$ $204.8$ $0.704$
Gong et al.[25] $13.15$ $0.119$ $199.6$ $0.682$
Peng et al.[62] $13.78$ $0.141$ $218.8$ $0.716$
Hayes et al.[28] $14.10$ $0.004$ $199.0$ $0.664$
Wang et al.[82] $13.63$ $0.0011$ $203.2$ $0.744$
Sparse-Standard $13.09$ $0.002$ $222.1$ $0.749$
SCA0.1 $\mathbf{12.89}$ $\mathbf{0.004}$ $\mathbf{228.5}$ $\mathbf{0.748}$
SCA0.25 $\mathbf{12.73}$ $\mathbf{0.004}$ $\mathbf{218.8}$ $\mathbf{0.737}$
SCA0.5 $\mathbf{12.42}$ $\mathbf{0.001}$ $\mathbf{231.9}$ $\mathbf{0.741}$
Medical No-Defense $31.48$ $0.935$ $10.66$ $0.998$
MNIST Gaussian-Noise $30.46$ $0.920$ $12.23$ $0.862$
GAN $27.34$ $0.480$ $33.77$ $0.998$
Gong et al.[25]++ $18.37$ $0.353$ $81.52$ $0.894$
Titcombe et al.[75] $21.33$ $0.431$ $30.60$ $0.899$
Gong et al.[25] $21.52$ $0.436$ $64.88$ $0.770$
Peng et al.[62] $19.05$ $0.420$ $107.9$ $0.908$
Hayes et al.[28] $18.48$ $0.007$ $150.9$ $0.824$
Wang et al.[82] $20.48$ $0.549$ $30.01$ $0.946$
Sparse-Standard $14.79$ $0.119$ $250.6$ $0.907$
SCA0.1 $\mathbf{13.43}$ $\mathbf{0.004}$ $\mathbf{352.1}$ $\mathbf{0.888}$
SCA0.25 $\mathbf{12.32}$ $\mathbf{0.004}$ $\mathbf{375.9}$ $\mathbf{0.882}$
SCA0.5 $\mathbf{12.04}$ $\mathbf{0.003}$ $\mathbf{369.9}$ $\mathbf{0.881}$
@@
Metric evaluations. Table 2 reports performance in the end-to-end network setting. In this setting, SCA’s performance advantage widens. In terms of PSNR & FID, training data reconstructions under SCA0.1 are degraded by factors of $1.1$ to $3.3$ and $1.1$ to $11.7$ compared to Peng et al. [62] and Wang et al.[82], respectively (but slightly worse SSIM vs. Wang et al. on CelebA). On all metrics, SCA0.1 outperforms Hayes et al.[28] and Titcombe et al. [75] by factors of $1.1$ to $2.4$ and $1.1$ to $107.7$, respectively. SCA0.1 also outperforms Gong et al. [25] by factors of $1.01$ to $109$. Increasing SCA’s $\lambda$ to $0.5$ causes it to outperform the same baselines on all metrics by factors of $1.2$ to $141$ [62], $1.2$ to $183$ [82], $1.2$ to $4.0$[28], $1.2$ to $191$ [75], and $1.1$ to $145.3$ [25], respectively.
@@
SCA’s performance is also as stable or more stable than baselines’ performance over multiple runs. For example, Fig. 5(a) plots the means & std. devs. of SCA & baselines’ per-run FID (the standard metric for faces) over multiple runs for CelebA faces in the Plug-&-Play setting, and Fig. 5(b) plots this for PSNR (the standard metric for grayscale images) on Medical MNIST in the end-to-end setting. SCA obtains better performance while also exhibiting stability of defense performance on par with the best baseline (and better stability than most baselines). See Appendix E.
@@
CelebA No-Defense $16.49$ $0.302$ $185.8$ $0.766$
Gaussian-Noise $15.44$ $0.227$ $191.1$ $0.753$
GAN $15.57$ $0.253$ $176.7$ $0.646$
Titcombe et al.[75] $14.99$ $0.144$ $194.2$ $0.725$
Gong et al.[25]++ $15.06$ $0.038$ $190.5$ $0.756$
Gong et al.[25] $15.65$ $0.044$ $185.8$ $0.653$
Peng et al.[62] $16.23$ $0.211$ $198.6$ $0.717$
Hayes et al.[28] $15.06$ $0.005$ $178.8$ $0.672$
Wang et al.[82] $14.82$ $0.173$ $189.6$ $0.652$
Sparse-Standard $15.39$ $0.009$ $187.0$ $0.746$
SCA0.1 $\mathbf{15.05}$ $\mathbf{0.005}$ $\mathbf{178.7}$ $\mathbf{0.745}$
SCA0.25 $\mathbf{14.76}$ $\mathbf{0.003}$ $\mathbf{191.1}$ $\mathbf{0.743}$
SCA0.5 $\mathbf{14.71}$ $\mathbf{0.003}$ $\mathbf{206.1}$ $\mathbf{0.739}$
Medical No-Defense $23.47$ $0.776$ $45.57$ $0.993$
MNIST Gaussian-Noise $21.93$ $0.722$ $44.72$ $0.811$
GAN $21.67$ $0.719$ $48.49$ $0.912$
Gong et al.[25]++ $21.07$ $0.573$ $67.53$ $0.931$
Titcombe et al.[75] $21.35$ $0.704$ $48.82$ $0.961$
Gong et al.[25] $21.33$ $0.720$ $41.74$ $0.925$
Peng et al.[62] $18.98$ $0.426$ $124.8$ $0.914$
Hayes et al.[28] $21.46$ $0.442$ $137.4$ $0.850$
Wang et al.[82] $20.03$ $0.538$ $65.17$ $0.986$
Sparse-Standard $15.33$ $0.149$ $142.4$ $0.955$
SCA0.1 $\mathbf{13.95}$ $\mathbf{0.008}$ $\mathbf{244.9}$ $\mathbf{0.946}$
SCA0.25 $\mathbf{12.31}$ $\mathbf{0.008}$ $\mathbf{255.3}$ $\mathbf{0.928}$
SCA0.5 $\mathbf{12.27}$ $\mathbf{0.001}$ $\mathbf{285.3}$ $\mathbf{0.909}$
@@
We also try varying SCA’s and Sparse-Standard’s sparsity parameters $\lambda$ and recompute PSNR, SSIM, FID, and accuracy. Appendix A.6 shows that for each $\lambda$ and defense metric, SCA significantly outperforms the off-the-shelf Sparse-Standard architecture for a small accuracy cost. Thus, for a given $\lambda$ with Sparse-Standard, we can use a (smaller) $\lambda$ with SCA to obtain better reconstruction and higher or equal (within $0.0017$) accuracy. SCA is also amenable to more sophisticated tuning (and performance improvements) by tuning different $\lambda$ per sparse layer (e.g., by having a sparser representation of inputs but less sparse reductions of downstream layers). We avoid such tuning here as it is unnecessary for good performance.
@@ 5 Empirical analysis of sparse coding robustness to attack
Sparse-coding layers’ robustness to privacy attacks can be observed empirically. Consider that the attacker trains the attack to map leaked raw hidden layer outputs back to input images. Attacks are thus highly dependent on these outputs’ distributions. Recall that UMAP projections compute a 2D visualization of the global structure of distances between different training images’ features according to a particular layer [50]. Fig. 6 plots UMAP 2D projections of linear layer feature distributions of training inputs after either two linear layers (Figs. 6(a) & 6(d)), two convolutional layers (Figs. 6(b) & 6(e)), or two sparse coding layers (with interspersed dense layers – Figs. 6(c) & 6(f)). Importantly, observe that after two linear or two convolutional layers, points are clustered by color, i.e., input images’ features are highly clustered by label. This class-clustered property leaves such layers vulnerable to model inversion attacks, as an attacker can ‘home in on’ examples from a specific class. In contrast, the goal in sparse coding is not to optimize the classification objective by separating classes, but rather to jettison unnecessary information. Here, this means that unnecessary information is jettisoned both from the input image and also the downstream dense layer. Per Figs. 6(c) & 6(f), this tends to ‘uncluster’ remaining non-sparsified features of training examples from the same class, making it much harder for an attacker to compute informative gradients to home in on a training example.
@@ 6 Discussion & Conclusion
In this paper, we have provided the first study of sparse coding-based neural network architectures that are robust to model inversion attacks. Specifically, we have shown that the natural properties of sparse coded layers can control the extraneous private information about the training data that is encoded in a network without resorting to complex and computationally intensive parameter tuning techniques. Our work reveals a deep connection between state-of-the-art privacy vulnerabilities and three decades of computer science research on sparse coding for other application domains. Currently, our basic research implementation of SCA achieves compute times that in the worst-case are no better than some SOTA baselines (see Appendix F). However, given the rich theoretic body of work on fast algorithms and provable guarantees for sparse coding, we believe these aspects are opportune areas for future improvements.
Acknowledgements
We are grateful for generous support from OpenAI, as well as the Dartmouth College Cybersecurity Cluster Research. This work was partially funded by the Center for Nonlinear Studies and the Information Science and Technology Institute’s Cyber Security Summer School at Los Alamos National Laboratory, as well as an award from the Department of Energy’s Advanced Scientific Computing Research program (#77902).
@@ Appendix 0.A Appendix
This is the supplementary document containing the additional results and details of our proposed Sparse Coding Architecture (SCA) formulations, as well as cluster details and additional preliminaries.
@@ 0.A.1 Reproducibility
In order to promote further research and standardize the evaluations of new defenses, we provide full cluster-ready PyTorch [61] implementations of SCA and all benchmarks as well as replication codes for all experiments on our project page at: https://sayantondibbo.github.io/SCA.
@@ 0.A.2 Adapting Rozell LCA to Convolutional Networks
Although the original LCA formulation [64] was introduced for the non-convolutional case, it is based on the general principle of feature-similarity-based competition between neurons within the same layer, which can be adapted to the convolutional setting via only two minimal changes to Equation 3 [74, 43]. In Rozell’s original formulation, $\Psi(t)$ can simply be recast from a matrix multiplication to a convolution between the input and dictionary. Second, the lateral interaction tensor, $\mathcal{G}$ in Equation 3, can also be recast from a matrix multiplication to a convolution between the dictionary and its transpose. Neuron membrane potential works as follows:
where $\tau$ is a time constant, $\Psi(t)=\mathcal{X}*\Omega$ is the neuron’s bottom-up drive from the input computed by taking the convolution, *, between the input, $\mathcal{X}$, and the dictionary, $\Omega$, and $-\mathcal{P}(t)$ is the leak term [74, 43].